OpenStudy (anonymous):

Andre received $60 for his birthday and put it in his piggy bank. Each week, he puts 3 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 3w + 60 where w is the number of weeks he has saved 3 dollars. His sister Lana makes a table of reasonable values for w. Lana selects 12 as a value for w. What is the corresponding table entry for D?

OpenStudy (anonymous):

well, try putting 12 as the value for w in that equation. I get D = 3 x 12 + 60 D = 36 + 60 D = 96 do you agree?

OpenStudy (anonymous):

Ah yes thank you <33